OpenStudy (anonymous):

10p-2(3p-6)=4(3p-6)-8p

OpenStudy (anonymous):

\[10p-2(3p-6)=4(3p-6)-8p\]distribute the coefficients into the parenthesis:\[10p-6p-12=12p-24-8p\]combine the like terms:\[4p-12=4p-24\]add 12 to both sides:\[4p=4p-12\]subtract 4p from both sides:\[0=-12\]since you have arrived at a false conclusion...\[\text{= no solution}\]

OpenStudy (anonymous):

hope thats helpful! @ashlynfc :)

OpenStudy (anonymous):

oh man, thank you. i've been stuck on this one. i keep going up to 36 instead to 12... appreciate the help.